OpenStudy (anonymous):

QRST is a rectangle. If RT = 3x - 8 and SQ = x + 16, find the value of x.

OpenStudy (er.mohd.amir):

these are diagonal of rectangle draw fig. and=x+ diagonal of rectangle are equal hence RT=SQ 3x-8=x+16 3x-x=16+8 2x=24 x=24/2 x=12 unit